# Permuton limits for some permutations avoiding a single pattern

### v2

@@ 1.1. Definitions
We note that classes of the form $\textbf{Av}_{n}(I_{k_{1}}\oplus J_{k_{2}}\oplus I_{k_{3}})$ are not necessarily in bijection with $\textbf{Av}_{n}(I_{k_{1}+k_{2}+k_{3}})$ and are much harder to analyze. In particular, this collection of classes contains $\textbf{Av}_{n}(1324)$ which is notoriously difficult to study. We suspect that the permuton limit associated to uniform sampling from these classes would also converge to $\mu_{J}$.
@@
If $A$ forms a decreasing sequence, the set $\text{SW}(A)$ will be called a Young diagram. We often view a permutation $\sigma$ of size $n$ as a subset of $R(n,n)$ by filling the squares $R((i-1,\sigma(i)-1),(i,\sigma(i))$ within the rectangle $R(n,n)$. For $\epsilon>0$ and $n\in\mathbb{N}$ we define three subregions of $R(n,n)$:
@@
For the bijection sending $\sigma$ to $\rho$ we say there is a frozen region in $R(n,n)$ where $\rho(i)=j$ if and only if $\sigma(i)=j$ for squares $(i,j)$ in this region. If $X$ is contained in this frozen region then $\mu_{\sigma}\left(\frac{1}{n}X\right)=\mu_{\rho}\left(\frac{1}{n}X\right)$. The bijection from $\rho^{rc}$ to $\pi^{rc}$ also has a similar frozen region.
@@
In Section 4 we explore the results in [6], where it is proven, among other results, that the points of $\sigma$ can be partitioned into a collection of $d=k_{1}+k_{2}+k_{3}-1$ decreasing sequences which are not too far from the line $x+y=n+1$. We establish Proposition 4.1, which says that with probability tending to 1 as $n$ increases, permutations $\sigma\in\textbf{Av}_{n}(I_{d})$ satisfy a set of good properties which then imply certain properties for permutations $\rho$, $\rho^{rc}$, $\pi^{rc}$ and $\pi$. In particular, for some fixed $\delta$ and $\delta^{\prime}$, if $\sigma$ is good, the frozen region under the bijection from $\sigma$ to $\rho$ and the frozen region under the bijection from $\rho^{rc}$ to $\pi^{rc}$ contain $W^{+}_{\delta,n}$ and $W^{+}_{\delta^{\prime},n}$, respectively. Moreover we have that $\mu_{\sigma}(W^{+}_{\delta})<\delta$ and $\mu_{\rho^{rc}}(W^{+}_{\delta^{\prime}})<\delta^{\prime}$, which combined with Lemma 3.1 eventually shows that $\mu_{\pi}(W_{\epsilon})<\epsilon$ for some $\epsilon$ which can be made arbitrarily small if $\delta$ and $\delta^{\prime}$ are sufficiently small.
For a sequence of permutations $\{\pi_{n}\}$ with $\pi_{n}\in\textbf{Av}_{n}(J_{k_{1}}\oplus I_{k_{2}}\oplus J_{k_{3}})$ chosen uniformly at random, the permuton $\mu_{\pi_{n}}$ will have $\mathbb{P}_{n}(\mu_{\pi_{n}}(W_{\epsilon})<\epsilon)\to 1$. Thus we can conclude that $\mu_{\pi_{n}}\to_{d}\mu_{J}$ where the convergence is in distribution.
@@ 2. Backelin-West-Xin Bijection
In this section, we discuss the bijection from [7], with some adjustments to their notation as shown in [2]. First, we define the notions of Wilf-equivalence and shape-Wilf equivalence. Let $S_{n}(\sigma)$ denote the set of permutations of size $n$ that do not contain an occurrence of a permutation $\sigma$. We say that two permutations $\sigma$ and $\sigma^{\prime}$ are Wilf-equivalent, denoted $\sigma\sim_{W}\sigma^{\prime}$, if $|S_{n}(\sigma)|=|S_{n}(\sigma^{\prime})|$. Hence, under this definition all length 3-avoiding permutations are Wilf equivalent, since for all $\tau\in S_{3}$, we have $|S_{n}(\tau)|=C_{n}$, the $n$th Catalan number.
@@
For shape-Wilf equivalence, consider a Young diagram $\lambda$, where $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$. We use the French convention for Young diagrams so the $i$th row starting from the bottom row has length $\lambda_{i}$. We define $L$ a traversal of $\lambda$ as a filling of boxes of $\lambda$ with zeros and ones, with exactly one 1 in each row and column; for our purposes, we will illustrate ones with dots, corresponding to entries of a permutation, and zeros with empty boxes. Given $\sigma\in S_{m}$, let $M_{\sigma}$ denote the traversal given by $\sigma$ of the $m\times m$ Young diagram. We say that $L$ contains a copy of $\sigma$ if $M_{\sigma}$ (including all its $1$s and 0s) is present in $L$ when restricting to rows $r_{1}<r_{2}<\dots<r_{m}$ and columns $c_{1}<c_{2}<\dots<c_{m}$. Hence, we view the pattern as being present only if it is contained in a complete rectangle contained within the Young diagram (See Figure 1). On the other hand, we say that $L$ avoids $\sigma$ if $L$ contains no copy of $\sigma$. Let $S_{\lambda}(\sigma)$ denote the traversals of $\lambda$ that avoid $\sigma$. Two permutations $\sigma$ and $\sigma^{\prime}$ are shape-Wilf equivalent, denoted $\sigma\sim_{sW}\sigma^{\prime}$, if $|S_{\lambda}(\sigma)|=|S_{\lambda}(\sigma^{\prime})|$ for all Young diagrams $\lambda$. For example, the shape-Wilf equivalence classes for length 3 permutations are $\{(213,132\}$, $\{123,231,321\}$, and $\{312\}$.
@@
Applying the Backelin-West-Xin bijection, for the example shown in Figure 2, the permutation avoids the pattern $21$ (by the previous discussion of pattern containment for Young diagrams).
@@ 3. Analytical Tools
The following proposition will be a useful tool in showing permuton collapse to the anti-diagonal.
@@
Fix small positive $\delta\in(0,.5)$. If $\mu$ is a permuton such that $\mu(W^{+}_{\delta})<\delta$, then for $\epsilon>2/\log(\delta^{-1})$, $\mu(W^{-}_{\epsilon})<\epsilon$ and $\mu(W_{2\epsilon})<2\epsilon$.
@@
The set $W^{-}_{\epsilon}$ is a subset $\bigcup_{0\leq i\leq m-2}H_{i}^{-}.$ Thus the proof will follow by showing that
@@ 4. Structure of permutations
For a permutation $\sigma\in\textbf{Av}_{n}(I_{d+1})$ of size $n$, and a subset of the indices $C\subset[n]$, we say $i\in C$ is a left-to-right minima of $C$ with respect to $\sigma$ if for $j\in C$ with $j<i$, $\sigma(j)>\sigma(i)$. Let $A^{1}_{\sigma}\subset[n]$ denote the left-to-right minima of $[n]$ with respect to $\sigma$. Then define recursively for $l>1$, $A^{l}_{\sigma}$ to be the left-to-right minima of $n]\setminus\bigcup_{i=1}^{l-1}A^{i}$ with respect to $\sigma$. This partitions $[n]$ into $d$ disjoint (possibly empty) subsets. This partition has the property that for $l>1$ and $i\in A^{l}_{\sigma}$, there exists a sequence of indices $i_{1}<\cdots<i_{l}=i$ such that $i_{j}\in A^{j}$ for $1\leq j\leq l$ and $\sigma(i_{1})<\cdots<\sigma(i_{l})$. This canonical partition does not guarantee the existence of a complete increasing sequence $i_{1}<\cdots<i_{l}=i<\cdots<i_{d}$ with $i_{j}\in A^{j}$ for $1\leq j\leq d$. However, we will show that for most permutations and most indices in $[n]$ such a complete increasing sequence does exist.
@@
for each $l\in[d]$. We translate results from [6] for $\sigma^{*}\in\textbf{Av}_{n}(J_{d})$ to our situation via the bijection $\sigma=\sigma^{*c}$. Translating Theorem 2.1 in [6] shows that this collection of functions converges in distribution to $\Lambda=(\lambda^{1},\cdots,\lambda^{d})$ where the $\lambda^{l}$’s are each a Brownian bridge from $(0,0)$ to $(1,0)$ jointly conditioned to satisfy:
@@
To show the last property holds for $\textbf{Av}_{n}(I_{d+1},\epsilon)$, we will assume the first four already hold. Suppose there exists $c>0$ such that for all $n>0$, the probability that there exists $1\leq l\leq d$ and some $i\in A^{l}_{\sigma}\bigcap[\epsilon n,(1-\epsilon)n]$ such that $0<\sigma(j^{l+1}(i))-\sigma(i)<n^{.4}$ is uniformly bounded below by $c$. Let $i=nt$ for some $t\in(\epsilon,1-\epsilon)$, and let $j^{l+1}(i)=nt-sn^{.2}$. By the second to last property of $\textbf{Av}_{n,\epsilon}(I_{d+1})$, $0<s<1$. Since $P_{\sigma}$ converges in distribution to $\Lambda$, we have that
$f^{l+1}_{\sigma}(t-sn^{-.8})-f^{l}_{\sigma}(t)\to\lambda^{l+1}(t)-\lambda^{l}(t).$
The left hand side of this equation is positive and bounded above by $n^{.4}/\sqrt{2dn}$ and thus tends to 0. However, the right hand side is bounded away from zero for $t\neq 0$ or $1.$ Thus, from this contradiction, we may suppose that $\sigma(j^{l+1}(i))-\sigma(i)>n^{.4}$ with probability tending to 1 as desired. ∎
Proposition 4.1 implies a useful lemma that will allow us to pin down the frozen region under certain Backelin-West-Xin bijections.
@@
Let $0<\epsilon_{1}$ and let $\epsilon_{2}>\max(4\epsilon_{1},2/\log(2\epsilon_{1})^{-1})$, $\epsilon>4/\log(\epsilon_{2}^{-1})$. For integers $k_{1}>0$, $k_{2}\geq 0$, and $k_{3}>0.$ Let $\pi\in\textbf{Av}_{n}(J_{k_{1}}\oplus I_{k_{2}}\oplus J_{k_{3}})$ be the image of $\sigma\in\textbf{Av}_{n,\epsilon_{1}}(I_{k_{1}}\oplus I_{k_{2}}\oplus I_{k_{3}}).$ Then $\mu_{\pi}(W_{\epsilon})<\epsilon.$
@@
Figure 4. A permutation, $\sigma\in\textbf{Av}_{n}(I_{5})$, and its image under the Backelin-West-Xin bijection, $\rho\in\textbf{Av}_{n}(J_{2}\oplus I_{3})$, with region $\text{SW}(A^{2}_{\sigma},\sigma)$ outlined by solid blue line. The frozen region under this bijection that is not shaded can properly contain the complement of $\text{SW}(A^{2}_{\sigma},\sigma)$ as the two lowest brown shaded points of $A^{1}_{\sigma}$ do not have a copy of $I_{3}$ north west of those squares.
Let $\rho$, $\rho^{rc}$, $\pi^{rc}$, and $\pi$ be the images under the appropriate bijections defined above. For any square not in $\text{SW}(A^{k_{1}}_{\sigma},\sigma)$, there can be no copy of $I_{k_{2}}\oplus I_{k_{3}}$ in $\sigma$ northwest of that square. Thus, the frozen region under the bijection from $\sigma$ to $\rho$ will contain the complement of $\text{SW}(A^{k_{1}}_{\sigma},\sigma).$ That is, for $(i,j)\notin\text{SW}(A^{k_{1}}_{\sigma},\sigma)$, $\rho(i)=j$ if and only if $\sigma(i)=j$. In particular, all points of the form $\{(i,\sigma(i)):i\in A^{l}_{\sigma}\}_{l\geq k_{1}}$ are contained in the frozen region. It is possible the frozen region may contain some points outside of this region’s complement as well, as shown in Figure 4. Note that by Proposition 4.1, the frozen region also contains the region
@@
Now consider the decreasing set of points $\{(i,\sigma(i)):i\in A^{k_{1}+k_{2}}_{\sigma}=\{(i,\rho(i)):i\in A^{k_{1}+k_{2}}_{\sigma}\}$. If $i^{\prime}\in A^{k_{1}+k_{2}}_{\sigma}$ is in the interval $(2\epsilon_{1}n,(1-2\epsilon_{1}n)$, then from Lemma 4.2, there exists an increasing sequence of indices $i^{\prime}=i_{k_{1}+k_{2}}<i_{k_{1}+k_{2}+1}<\dots<i_{k_{1}+k_{2}+k_{3}-1}$ with $i_{l}\in A^{l}_{\sigma}$ such that $(i_{l},\sigma(i_{l}))=(i_{l},\rho(i_{l}))$ is increasing for $k_{1}+k_{2}\leq l<k_{1}+k_{2}+k_{3}.$ Let $A=A^{k_{1}+k_{2}}\cap(2\epsilon_{1}n,(1-2\epsilon)n.$ Since $\rho$ avoids $J_{k_{1}}\oplus I_{k_{2}}\oplus I_{k_{3}}$, then $\rho$ will avoid $J_{k_{1}}\oplus I_{k_{2}}$ in the region $\text{SW}(A,\rho)$.
@@
Fix $\epsilon>0$. Choose $\epsilon_{1}$ small enough such that Lemma 5.1 applies. If $\sigma_{n}\in\textbf{Av}_{n,\epsilon_{1}}(I_{k_{1}}\oplus I_{k_{2}}\oplus I_{k_{3}})$, then $\pi_{n}\in\textbf{Av}_{n}(J_{k_{1}}\oplus I_{k_{2}}\oplus J_{k_{3}})$ satisfies $\mu_{\pi_{n}}(W_{\epsilon})<\epsilon.$ Let $\pi_{n}$ be chosen uniformly at random from the image of $\textbf{Av}_{n,\epsilon_{1}}(I_{k_{1}}\oplus I_{k_{2}}\oplus I_{k_{3}})\subset\textbf{Av}_{n}(J_{k_{1}}\oplus I_{k_{2}}\oplus J_{k_{3}})$. Then by Lemma 5.1, $\mu_{\pi_{n}}(W_{\epsilon})<\epsilon$, and thus by Lemma 3.2, $\mu_{\pi_{n}}\to\mu_{J}$ weakly. By Proposition 4.1, $\pi_{n}$ is the in the image of $\textbf{Av}_{n,\epsilon_{1}}(I_{k_{1}}\oplus I_{k_{2}}\oplus I_{k_{3}})$ with probability tending to 1 as $n$ increases, thus we may conclude that $\mu_{\pi_{n}}$ converges in distribution to $\mu_{J}$ with respect to the appropriate topology. ∎
Open Questions
Our results only look at some classes of permutation that are in bijection with those avoiding a monotone increasing sequence. However, Lemma 3.1 could be useful for showing permuton limits of $\mu_{J}$ for a larger set of pattern-avoiding classes. It would be nice to understand more about the permuton limits for classes avoiding a single pattern of length 4. Initially we attempted to explore $\textbf{Av}_{n}(1342)$ using the bijection developed by Bóna [3], though we were unable to find the permuton limit. Understanding permuton limits for $\textbf{Av}_{n}(1342)$, $\textbf{Av}_{n}(1324)$ and $\textbf{Av}_{n}(2413)$ would complete the picture for permuton limits of classes avoiding a single pattern of size $4$. We conjecture that both $\textbf{Av}_{n}(1342)$ and $\textbf{Av}_{n}(1324)$ have permuton limits of $\mu_{J}$, whereas for $\textbf{Av}_{n}(2413)$ the permuton limit is a bit of a mystery still. Observing that permutations in $\textbf{Av}_{n}(2413)$ have rotational symmetry shows that neither $\mu_{J}$ nor $\mu_{I}$ could be the permuton limit for this case. The permutation class $\textbf{Av}_{n}(2413,3142)$ was shown to have a random permuton limit [1] which suggests the permuton limits for the larger classes ($\textbf{Av}_{n}(2413)$ or $\textbf{Av}_{n}(3142)$) might have more complicated structure than what we explore in this paper.

### v3

@@ 1.1. Definitions
This theorem provides partial progress to the conjecture found on page 282 of [8] which states that any permutation class of the form $\textbf{Av}_{n}(\alpha\oplus\beta)$ will have a permuton limit of $\mu_{J}$. Our proof relies strongly on a bijection with $\textbf{Av}_{n}(I_{k_{1}+k_{2}+k_{3}})$. We note that classes of the form $\textbf{Av}_{n}(I_{k_{1}}\oplus J_{k_{2}}\oplus I_{k_{3}})$ are not necessarily in bijection with $\textbf{Av}_{n}(I_{k_{1}+k_{2}+k_{3}})$ and are much harder to analyze. In particular, this collection of classes contains $\textbf{Av}_{n}(1324)$ which is notoriously difficult to study.
@@
If $i_{k}$ is increasing, and $j_{k}$ is decreasing, the set $\text{SW}(A)$ will be called a Young diagram. We often view a permutation $\sigma$ of size $n$ as a subset of $R(n,n)$ by filling the boxes $R((i-1,\sigma(i)-1),(i,\sigma(i))$ within the rectangle $R(n,n)$. For $\epsilon>0$ and $n\in\mathbb{N}$ we define three subregions of $R(n,n)$:
@@
For the bijection sending $\sigma$ to $\rho$ we say there is a frozen region in $R(n,n)$ where $\rho(i)=j$ if and only if $\sigma(i)=j$ for boxes $(i,j)$ in this region. If $X$ is contained in this frozen region then $\mu_{\sigma}\left(\frac{1}{n}X\right)=\mu_{\rho}\left(\frac{1}{n}X\right)$. The bijection from $\rho^{rc}$ to $\pi^{rc}$ also has a similar frozen region.
@@
In Section 4 we explore the results in [6], where it is proven, among other results, that the points of $\sigma$ can be partitioned into a collection of $d=k_{1}+k_{2}+k_{3}-1$ decreasing sequences which are not too far from the line $x+y=n+1$ and which satisfy certain use properties with high probability. We make explicit these properties by defining a set $\textbf{Av}_{n,\epsilon}(I_{d})$ and show that $\sigma$ chosen uniformly from $\textbf{Av}_{n}(I_{d})$ lies in $\textbf{Av}_{n,\epsilon}(I_{d})$ with probability tending to 1. Lemmas 4.2 combines with 5.1 to show that if $\epsilon_{1}<\epsilon_{2}<\epsilon$ satisfy certain conditions and $\sigma\in\textbf{Av}_{n,\epsilon_{1}}(I_{d})$, the image $\pi\in\textbf{Av}_{n}(J_{k_{1}}\oplus I_{k_{2}}\oplus J_{k_{3}})$ with have a permuton that satisfies $\mu_{\pi}(W_{\epsilon})<\epsilon)$.
This is seen by first showing the frozen region under the bijection from $\sigma$ to $\rho$ contains the region $W^{+}_{n,2\epsilon_{1}}$ and the corresponding permutons satisfy $\mu_{\sigma}(W^{+}_{2\epsilon_{1}})=\mu_{\rho}(W^{+}_{2\epsilon_{1}})=0$ and thus $\mu_{\rho}(W_{\epsilon_{2}})<\epsilon_{2}$. Then through a slightly more complicated argument we show that the frozen region under the bijection from $\rho^{rc}$ to $\pi^{rc}$ contains the region $W^{+}_{n,2\epsilon_{2}}$ with $\mu_{\rho^{rc}}(W^{+}_{2\epsilon_{2}})=\mu_{\pi^{rc}}(W^{+}_{2\epsilon_{2}})<2\epsilon_{2}.$ Another application of Lemma 3.1 arrives at the above conclusion.
One difficultly of the proof is finding a way to show that the frozen region under the bijection that maps $\rho^{rc}$ to $\pi^{rc}$ is sufficiently large enough to contain a region of the form $W^{+}_{n,2\epsilon_{2}}$. If the frozen region under this bijection were too small, our proof would fall apart. However, using results developed in [6], Lemmas 4.1 and 4.2 provides a set of conditions that hold with high probability. The proof of Lemma 5.1 uses those properties to show that frozen region will, in fact, be large enough.
We can then conclude that for a sequence of permutations $\{\pi_{n}\}$ with $\pi_{n}\in\textbf{Av}_{n}(J_{k_{1}}\oplus I_{k_{2}}\oplus J_{k_{3}})$ chosen uniformly at random, the permuton $\mu_{\pi_{n}}$ will have $\mathbb{P}_{n}(\mu_{\pi_{n}}(W_{\epsilon})<\epsilon)\to 1$, and thus $\mu_{\pi_{n}}\to_{d}\mu_{J}$ where the convergence is in distribution with respect to topology of weak convergence.
@@ 2. Backelin-West-Xin Bijection
In this section, we discuss the bijection from [7], with some adjustments to their notation as shown in [2]. First, we define the notions of Wilf-equivalence and shape-Wilf equivalence. Let $S_{n}(\sigma)$ denote the set of permutations of size $n$ that do not contain an occurrence of a permutation $\sigma$. We say that two permutations $\sigma$ and $\sigma^{\prime}$ are Wilf-equivalent, denoted $\sigma\sim_{W}\sigma^{\prime}$, if $|S_{n}(\sigma)|=|S_{n}(\sigma^{\prime})|$ for all $n$. Hence, under this definition all length 3-avoiding permutations are Wilf equivalent, since for all $\tau\in S_{3}$, we have $|S_{n}(\tau)|=C_{n}$, the $n$th Catalan number.
@@
For shape-Wilf equivalence, consider a Young diagram $\lambda$, where $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$. We use the French convention for Young diagrams so the $i$th row starting from the bottom row has length $\lambda_{i}$. We define $L$ a traversal of $\lambda$ as a filling of boxes of $\lambda$ with zeros and ones, with exactly one 1 in each row and column; for our purposes, we will illustrate ones with dots, corresponding to entries of a permutation, and zeros with empty boxes. Given $\sigma\in S_{m}$, let $M_{\sigma}$ denote the traversal given by $\sigma$ of the $m\times m$ Young diagram. We say that $L$ contains a copy of $\sigma$ if $M_{\sigma}$ (including all its $1$s and 0s) is present in $L$ when restricting to rows $r_{1}<r_{2}<\dots<r_{m}$ and columns $c_{1}<c_{2}<\dots<c_{m}$. Hence, we view the pattern as being present only if it is contained in a complete rectangle contained within the Young diagram (See Figure 1). On the other hand, we say that $L$ avoids $\sigma$ if $L$ contains no copy of $\sigma$. Let $S_{\lambda}(\sigma)$ denote the traversals of $\lambda$ that avoid $\sigma$. Two permutations $\sigma$ and $\sigma^{\prime}$ are shape-Wilf equivalent, denoted $\sigma\sim_{sW}\sigma^{\prime}$, if $|S_{\lambda}(\sigma)|=|S_{\lambda}(\sigma^{\prime})|$ for all Young diagrams $\lambda$. For example, the shape-Wilf equivalence classes for length 3 permutations are $\{213,132\}$, $\{123,231,321\}$, and $\{312\}$.
@@
Applying the Backelin-West-Xin bijection, for the example shown in Figure 2, the traversal $\pi^{\prime}_{\lambda}$ avoids the pattern $21$ (by the previous discussion of pattern containment for Young diagrams).
@@ 3. Analytical Tools
The following lemma will be a useful tool in showing permuton collapse to the anti-diagonal.
@@
Fix small positive $\delta\in(0,.5)$. If $\mu$ is a permuton such that $\mu(W^{+}_{\delta})<\delta$, then for $\epsilon>2/\log(\delta^{-1})>\delta$, $\mu(W^{-}_{\epsilon})<\epsilon$ and $\mu(W_{2\epsilon})<2\epsilon$.
@@
The set $W^{-}_{\epsilon}$ is a subset of $\bigcup_{0\leq i\leq m-2}H_{i}^{-}.$ Thus the proof will follow by showing that
@@ 4. Structure of permutations
For a permutation $\sigma\in\textbf{Av}_{n}(I_{d+1})$ of size $n$, and a subset of the indices $C\subset[n]$, we say $i\in C$ is a left-to-right minimum of $C$ with respect to $\sigma$ if for $j\in C$ with $j<i$, $\sigma(j)>\sigma(i)$. Let $A^{1}_{\sigma}\subset[n]$ denote the left-to-right minima of $[n]$ with respect to $\sigma$. Then define recursively for $l>1$, $A^{l}_{\sigma}$ to be the left-to-right minima of $n]\setminus\bigcup_{i=1}^{l-1}A^{i}$ with respect to $\sigma$. This partitions $[n]$ into $d$ disjoint (possibly empty) subsets. This partition has the property that for $l>1$ and $i\in A^{l}_{\sigma}$, there exists a sequence of indices $i_{1}<\cdots<i_{l}=i$ such that $i_{j}\in A^{j}$ for $1\leq j\leq l$ and $\sigma(i_{1})<\cdots<\sigma(i_{l})$. This canonical partition does not guarantee the existence of a complete increasing sequence $i_{1}<\cdots<i_{l}=i<\cdots<i_{d}$ with $i_{j}\in A^{j}$ for $1\leq j\leq d$. However, we will show that for most permutations and most indices in $[n]$ such a complete increasing sequence does exist.
@@
for each $l\in[d]$. We translate results from [6] for $\sigma^{*}\in\textbf{Av}_{n}(J_{d})$ to our situation via the bijection $\sigma=\sigma^{*c}$. Translating Theorem 2.1 in [6] shows that this collection of functions converges in distribution with respect to the sup norm topology to $\Lambda=(\lambda^{1},\cdots,\lambda^{d})$ where the $\lambda^{l}$’s are each a Brownian bridge from $(0,0)$ to $(1,0)$ jointly conditioned to satisfy:
@@
To show the last property holds for $\textbf{Av}_{n}(I_{d+1},\epsilon)$, we will assume the first four are already true. Let $X$ be the $\inf_{1\leq l<d,t\in[\epsilon,1-\epsilon]}(\lambda^{l+1}(t)-\lambda^{l}(t))$. Let $Y_{n}$ be similar defined for $(f^{1}_{\sigma},\cdots,f^{d}_{\sigma})$ where $\sigma$ is chosen uniformly form $\textbf{Av}_{n}(J_{d}).$ The second property in the definition of Dyson Brownian bridge along with continuity of the component paths combine to show that $\mathbb{P}(X>c)\to 1$ as $c\to 0.$ Convergence in distribution of $(f^{1}_{\sigma},\cdots,f^{d}_{\sigma})$ to $(\lambda^{1},\cdots,\lambda^{d})$ implies that $Y_{n}\to_{d}X$ as well, and thus $\mathbb{P}(Y_{n}>n^{-.05})$ tends to 1 as $n$ increases. If $Y_{n}>n^{-.05}$, then $\sqrt{2dn}(f^{l+1}_{\sigma}(t)-f^{l}_{\sigma}(t))>n^{.45}$. Fix $t\in[\epsilon,1-\epsilon]$. Let $i=\lfloor nt\rfloor$. Then $i$ is in $A^{l}_{\sigma}$ for some $l$. If $l<d$, let $n(t-sn^{-.8})=j^{l+1}(i)$. By the second to last property of $\textbf{Av}_{n,\epsilon}(I_{d+1})$, $s$ will lie in the interval $(0,1)$. Using the third property we see that $\sqrt{2dn}(f^{l+1}_{\sigma}(t)-f^{l+1}_{\sigma}(t-sn^{-.8}))<n^{.2}.$ Combining everything we have with probability that tends to 1,
$\sigma(j^{l+1}(i))-\sigma(i)>$ $\sqrt{2dn}(f^{l+1}_{\sigma}((t-sn^{-.8})-f^{l}_{\sigma}(t))-n^{.2}$
$>$ $\sqrt{2dn}(f^{l+1}_{\sigma}(t)-f^{l}_{\sigma}(t))-2n^{.2}$
$>$ $n^{.45}-2n^{.2}$
$>$ $n^{.4}.$
Lemma 4.1 implies a useful lemma that will allow us to pin down the frozen region under certain Backelin-West-Xin bijections.
@@
Let $\epsilon_{1}>0$ and let $\epsilon_{2}>\max(4\epsilon_{1},2/\log(2\epsilon_{1})^{-1})$, $\epsilon>4/\log(\epsilon_{2}^{-1})$. For integers $k_{1}>0$, $k_{2}\geq 0$, and $k_{3}>0.$ Let $\pi\in\textbf{Av}_{n}(J_{k_{1}}\oplus I_{k_{2}}\oplus J_{k_{3}})$ be the image of $\sigma\in\textbf{Av}_{n,\epsilon_{1}}(I_{k_{1}}\oplus I_{k_{2}}\oplus I_{k_{3}}).$ Then $\mu_{\pi}(W_{\epsilon})<\epsilon.$
@@
Figure 4. A permutation, $\sigma\in\textbf{Av}_{n}(I_{5})$, and its image under the Backelin-West-Xin bijection, $\rho\in\textbf{Av}_{n}(J_{2}\oplus I_{3})$, with region $\text{SW}(A^{2}_{\sigma},\sigma)$ outlined by solid blue line. The frozen region under this bijection that is not shaded can properly contain the complement of $\text{SW}(A^{2}_{\sigma},\sigma)$ as the two lowest brown shaded points of $A^{1}_{\sigma}$ do not have a copy of $I_{3}$ northeast of those boxes.
Let $\rho$, $\rho^{rc}$, $\pi^{rc}$, and $\pi$ be the images under the appropriate bijections defined above. For any box not in $\text{SW}(A^{k_{1}}_{\sigma},\sigma)$, there can be no copy of $I_{k_{2}}\oplus I_{k_{3}}$ in $\sigma$ northwest of that box. Thus, the frozen region under the bijection from $\sigma$ to $\rho$ will contain the complement of $\text{SW}(A^{k_{1}}_{\sigma},\sigma).$ That is, for $(i,j)\notin\text{SW}(A^{k_{1}}_{\sigma},\sigma)$, $\rho(i)=j$ if and only if $\sigma(i)=j$. In particular, all points of the form $\{(i,\sigma(i)):i\in A^{l}_{\sigma}\}_{l\geq k_{1}}$ are contained in the frozen region. It is possible the frozen region may contain some points outside of this region’s complement as well, as shown in Figure 4. Note that by Lemma 4.1, the frozen region also contains the region
@@
Now consider the decreasing set of points $\{(i,\sigma(i)):i\in A^{k_{1}+k_{2}}_{\sigma}=\{(i,\rho(i)):i\in A^{k_{1}+k_{2}}_{\sigma}\}$. If $i^{\prime}\in A^{k_{1}+k_{2}}_{\sigma}$ is in the interval $(2\epsilon_{1}n,(1-2\epsilon_{1})n)$, then from Lemma 4.2, there exists an increasing sequence of indices $i^{\prime}=i_{k_{1}+k_{2}}<i_{k_{1}+k_{2}+1}<\dots<i_{k_{1}+k_{2}+k_{3}-1}$ with $i_{l}\in A^{l}_{\sigma}$ such that $(i_{l},\sigma(i_{l}))=(i_{l},\rho(i_{l}))$ is increasing for $k_{1}+k_{2}\leq l<k_{1}+k_{2}+k_{3}.$ Let $A=A^{k_{1}+k_{2}}\cap(2\epsilon_{1}n,(1-2\epsilon_{1})n).$ Since $\rho$ avoids $J_{k_{1}}\oplus I_{k_{2}}\oplus I_{k_{3}}$, then $\rho$ will avoid $J_{k_{1}}\oplus I_{k_{2}}$ in the region $\text{SW}(A,\rho)$.
@@
Fix $\epsilon>0$. Choose $\epsilon_{1}$ small enough such that Lemma 5.1 applies. If $\sigma_{n}\in\textbf{Av}_{n,\epsilon_{1}}(I_{k_{1}}\oplus I_{k_{2}}\oplus I_{k_{3}})$, then $\pi_{n}\in\textbf{Av}_{n}(J_{k_{1}}\oplus I_{k_{2}}\oplus J_{k_{3}})$ satisfies $\mu_{\pi_{n}}(W_{\epsilon})<\epsilon.$ Let $\pi_{n}$ be chosen uniformly at random from the image of $\textbf{Av}_{n,\epsilon_{1}}(I_{k_{1}}\oplus I_{k_{2}}\oplus I_{k_{3}})\subset\textbf{Av}_{n}(J_{k_{1}}\oplus I_{k_{2}}\oplus J_{k_{3}})$. Then by Lemma 5.1, $\mu_{\pi_{n}}(W_{\epsilon})<\epsilon$, and thus by Lemma 3.2, $\mu_{\pi_{n}}\to\mu_{J}$ weakly. By Lemma 4.1, $\pi_{n}$ is the in the image of $\textbf{Av}_{n,\epsilon_{1}}(I_{k_{1}}\oplus I_{k_{2}}\oplus I_{k_{3}})$ with probability tending to 1 as $n$ increases, thus we may conclude that $\mu_{\pi_{n}}$ converges in distribution to $\mu_{J}$ with respect to the appropriate topology. ∎
6. Open Questions
Our results only look at some classes of permutation that are in bijection with those avoiding a monotone increasing sequence. However, Lemma 3.1 could be useful for showing permuton limits of $\mu_{J}$ for a larger set of pattern-avoiding classes. Initially we attempted to explore $\textbf{Av}_{n}(1342)$ using results developed by Bóna [3]. We could not show a permuton limit, but suspect convergence to $\mu_{J}$ could be shown by understanding how the bijection in [3] with certain labeled trees could give a one sided bound on the permuton limit. It would be nice to understand more about the permuton limits for classes avoiding a single pattern of length 4. Understanding permuton limits for $\textbf{Av}_{n}(1342)$, $\textbf{Av}_{n}(1324)$ and $\textbf{Av}_{n}(2413)$ would complete the picture for permuton limits of classes avoiding a single pattern of size $4$. We conjecture that both $\textbf{Av}_{n}(1342)$ and $\textbf{Av}_{n}(1324)$ have permuton limits of $\mu_{J}$, whereas for $\textbf{Av}_{n}(2413)$ the permuton limit is a bit of a mystery still. Observing that permutations in $\textbf{Av}_{n}(2413)$ have rotational symmetry shows that neither $\mu_{J}$ nor $\mu_{I}$ could be the permuton limit for this case. The permutation class $\textbf{Av}_{n}(2413,3142)$ was shown to have a random permuton limit [1] which suggests the permuton limits for the larger classes ($\textbf{Av}_{n}(2413)$ or $\textbf{Av}_{n}(3142)$) might have more complicated structure than what we explore in this paper.